# Spatial Phonons: A Phenomenological Viscous Dark Energy Model for DESI

### v1

@@ 2.4 A Caveat for the Phonon Vacuum Energy Density
where both $T_{s}^{\text{geom}}$ and $F(b)$ are understood as renormalized quantities. We expand around a homogeneous equilibrium configuration $b=b_{0}+\delta b$ and write,
@@ 2.5 Effective Potential $F(b)$ Near Equilibrium
This example shows explicitly that the phenomenological parameters $\varepsilon$ and $\kappa$ which we will use in Secs. 3 and 4 can be viewed as the first two derivatives of a local potential $F(b)$ evaluated at the equilibrium configuration $b_{0}$. Higher order terms in (2.59) are not needed for the small homogeneous deformations that are relevant for the late time cosmological evolution considered in this work.
@@ 3.2 Physical picture on the Elastic Brane
In an accelerating universe there is a Gibbons-Hawking temperature associated de Sitter horizon [10],
$T_{H}\simeq\frac{H}{2\pi}$ (3.6)
where $H$ is the Hubble parameter. The phonon sector is therefore in contact with a thermal bath at temperature $T_{H}$. We assume that the phonon sector can be described within near equilibrium hydrodynamics with respect to this effective thermal bath. Concretely, we assume that the microscopic processes that create and annihilate phonons and that redistribute their momentum are fast enough that the phonon distribution relaxes toward a local equilibrium characterized by $T_{H}$ on time scales shorter than or comparable to the Hubble time during the epoch when bulk viscosity is relevant. Under this assumption the Kubo formula and the Maxwell viscoelastic model provide a meaningful characterization of the bulk viscosity $\zeta$ in terms of correlation functions evaluated at temperature $T_{H}$. We will discuss these ideas in the Sec. 3.3.
@@ 4 Cosmological Implications and Comparison with DESI
$w_{\rm CPL}(0)\simeq-0.856,$
$w_{\rm CPL}(0.5)\simeq-1.033,$
$w_{\rm CPL}(1.0)\simeq-1.121,$
$w_{\rm CPL}(1.1)\simeq-1.134,$
$w_{\rm CPL}(1.2)\simeq-1.145,$
$w_{\rm CPL}(1.3)\simeq-1.156,$
$w_{\rm CPL}(1.4)\simeq-1.165,$
$w_{\rm CPL}(1.5)\simeq-1.174$ (4.7)
@@
Using the above choice of parameters we obtain the effective equation of state values
$w_{\rm eff}(0)\approx-0.820,$
$w_{\rm eff}(0.5)\approx-0.971,$
$w_{\rm eff}(1.0)\approx-1.135,$
$w_{\rm eff}(1.1)\approx-1.153,$
$w_{\rm eff}(1.2)\approx-1.163,$
$w_{\rm eff}(1.3)\approx-1.167,$
$w_{\rm eff}(1.4)\approx-1.165,$
$w_{\rm eff}(1.5)\approx-1.158$ (4.10)
@@
It is worth noting that the DESI motivated CPL fit selects a parameter set that is quite natural from the point of view of the viscoelastic model. The values $\kappa\simeq\varepsilon\simeq 1/3$ are less than unity, as expected, and both represent a moderate fraction of the brane tension $T_{s}$. At the same time the associated phonon sound speed, $c_{s}^{2}=\kappa/\varepsilon\simeq 1$, lies very close to the causal limit without overshooting it. This is a plausible outcome for longitudinal modes in a very stiff medium and it shows that the DESI driven fit does not require an exotic or finely tuned region of parameter space. Instead fitting DESI using our viscous dark energy model naturally singles out a regime in which the brane behaves like a strongly elastic but still well behaved effective medium where the phonon enthalpy and bulk modulus are each of order one third of the geometric tension $T_{s}$ and the phonons are naturally causal. As such, given below are tables summarizing the findings.
@@
Finally, we now close by commenting on the mass scale of the phonons.
In the simplified description of Section 3 we model the relaxation time as a function of the Hubble rate and introduce a characteristic scale $H_{\star}$ at which the viscous response is strongest. We identify this scale with the phonon mass through,
@@ 5 Conclusion and Outlook
A second important step is to move beyond the homogeneous background and analyze cosmological perturbations in this framework. The phonon sound speed, the viscous pressure and any induced anisotropic stress will modify the growth of density perturbations and the evolution of metric potentials. One can therefore confront the model with large scale structure data, redshift space distortion measurements and lensing observations in addition to background probes. A consistent treatment of scalar perturbations would place quantitative bounds on $\kappa$, $\varepsilon$, the crossing scale $H_{\star}$ and the effective phonon mass, and it would reveal whether the viscoelastic dark energy sector remains stable and well behaved at the level of fluctuations.
@@ 6 Acknowledgment
The author acknowledges the use of ChatGPT-5.1 (October 2025 version) to edit and refine the tone and clarity of selected sections of this manuscript. The model was used solely to polish and rephrase text originally written by the author. All passages prepared with LLM assistance were reviewed, revised where needed and approved by the author, who retains full responsibility for the final content.

### v2

@@ 2.4 A Caveat for the Phonon Vacuum Energy Density
where both $T_{s}$ and $F(b)$ are understood as renormalized quantities. We expand around a homogeneous equilibrium configuration $b=b_{0}+\delta b$ and write,
@@ 2.5 Effective Potential $F(b)$ Near Equilibrium
This example shows explicitly that the phenomenological parameters $\varepsilon$ and $\kappa$ which we will use in Secs. 3 and 4 can be viewed as the first two derivatives of a local potential $F(b)$ evaluated at the equilibrium configuration $b_{0}$. Higher order terms in (2.59) are not needed for the small homogeneous deformations that are relevant for the late time cosmological evolution considered in this work. We keep $\varepsilon$ and $\kappa$ as phenomenological parameters as mentioned before. We not derive them here from first principles, reserving that analysis for a future work.
@@ 3.2 Physical picture on the Elastic Brane
In this work, we will interpret the Gibbons-Hawking temperature [10] of a spatially flat FRW spacetime,
$T_{H}\equiv\frac{H}{2\pi}$ (3.6)
as an effective horizon temperature associated with the cosmological apparent horizon, rather than as literal Hawking radiation from an exact de Sitter event horizon. This viewpoint follows the horizon-thermodynamics approach in which the Friedmann equations can be written as a first law $dE=T_{H}dS+WdV$ at the apparent horizon, with $T_{H}=1/2\pi R_{A}$ and $R_{A}=1/H$ for $k=0$ FRW backgrounds [30, 31, 32, 37]. We assume that the phonon sector is in near-equilibrium with this effective thermal bath at temperature $T_{H}$ during the epoch when bulk viscosity is relevant. Concretely, we assume that the microscopic processes that create and annihilate phonons and that redistribute their momentum are fast enough that the phonon distribution relaxes toward a local equilibrium characterized by $T_{H}$ on time scales shorter than or comparable to the Hubble time during the epoch when bulk viscosity is relevant. Under this assumption the Kubo formula and the Maxwell viscoelastic model provide a meaningful characterization of the bulk viscosity $\zeta$ in terms of correlation functions evaluated at temperature $T_{H}$. We will discuss these ideas in the Sec. 3.3.
@@ 4 Cosmological Implications and Comparison with DESI
$z$ $w_{\rm CPL}(z)$
0.0 0.856-0.856
0.5 1.033-1.033
1.0 1.121-1.121
1.1 1.134-1.134
1.2 1.145-1.145
1.3 1.156-1.156
1.4 1.165-1.165
1.5 1.174-1.174
@@
Using the above choice of parameters we obtain the effective equation of state values as,
$z$ $w_{\rm eff}(z)$
0.0 0.820-0.820
0.5 0.971-0.971
1.0 1.135-1.135
1.1 1.153-1.153
1.2 1.163-1.163
1.3 1.167-1.167
1.4 1.165-1.165
1.5 1.158-1.158
@@
It is worth noting that the DESI motivated CPL fit selects a parameter set that is quite natural from the point of view of the viscoelastic model. The values $\kappa\simeq\varepsilon\simeq 1/3$ are less than unity, as expected, and both represent a moderate fraction of the brane tension $T_{s}$. At the same time the associated phonon sound speed, $c_{s}^{2}=\kappa/\varepsilon\simeq 1$, lies very close to the causal limit without overshooting it. This is a plausible outcome for longitudinal modes in a very stiff medium and it shows that the DESI driven fit does not require an exotic or finely tuned region of parameter space. Instead fitting DESI using our viscous dark energy model naturally singles out a regime in which the brane behaves like a strongly elastic but still well behaved effective medium where the phonon enthalpy and bulk modulus are each of order one third of the geometric tension $T_{s}$ and the phonons are naturally causal. As such, given below are tables summarizing the findings and a comparison plot.
@@
Figure 1: Comparison between the Chevallier–Polarski–Linder equation of state $w_{\rm CPL}(z)$ (solid line) with parameters $(w_{0},w_{a})=(-0.856,-0.53)$ and the effective equation of state $w_{\rm eff}(z)$ (dashed line) of the viscoelastic spatial phonon model for $\varepsilon=0.335$, $\kappa=0.33489$ and $H_{\star}/H_{0}=2.10$. The horizontal dotted line shows the cosmological constant value w=1w=-1. Over the DESI redshift range $0\leq z\lesssim 1.6$, the viscous dark energy model closely tracks the CPL fit.
We now close by commenting on the mass scale of the phonons.
In the simplified description of Section 3 we modeled the relaxation time as a function of the Hubble rate and introduced a characteristic scale $H_{\star}$ at which the viscous response is strongest. We identified this scale with the phonon mass through,
@@ 5 Conclusion and Outlook
A second important step is to move beyond the homogeneous background and analyze cosmological perturbations in this framework. The phonon sound speed, the viscous pressure and any induced anisotropic stress could in principle affect the growth of density perturbations and the evolution of metric potentials. This makes it natural to confront the model with large scale structure observables such as redshift space distortion measurements and lensing data, in addition to background probes. At present it is not clear whether the perturbative effects of the viscoelastic sector would lead to observable deviations or remain hidden within current uncertainties. Only a detailed and consistent treatment of scalar perturbations can answer this question in a quantitative manner. Such an analysis would place bounds on $\kappa$, $\varepsilon$, the crossing scale $H_{\star}$ and the effective phonon mass, and would test whether the viscoelastic dark energy sector remains stable and well behaved at the level of fluctuations.
@@ 6 Acknowledgment
The author acknowledges the use of ChatGPT-5.1 (November 2025 version) to edit and refine the tone and clarity of selected sections of this manuscript. The model was used solely to polish and rephrase text originally written by the author. All passages prepared with LLM assistance were reviewed, revised where needed and approved by the author, who retains full responsibility for the final content.